Pressure in a container of 1,000 moles H2O with V = 14.70 L when T = 295.0 K. Use three decimals in the answer.
Answer #1

Assuming H2O obeys idea gas law,

We have in problem

Volume of container = 14.70 liters = (14.70 / 1000) m3 = 0.01470 m3

No of moles of H2O = 1000 moles = 1 Kmol

The temperature of container = 295 K

Idea gas law,

PV = nRT

Solving for P (Pressure)

P = nRT / V

where

R = Constant = 8.314 kJ / Kmol. K

Putting all values in above equation

P = (1 * 8.314 * 295) / 0.01470

P = 166,845.5782 KJ / m3 = 166,845.7582 K.Pa = 166.8457582 MPa